OpenStudy (anonymous):

Wayne had $982 and Oscar had $3517. Every month, Wayne saves $48 while Oscar saves $9. How many months will it take for both to have the same amount of money?

OpenStudy (anonymous):

let the number of months after which they have same amount of money = x after x months, money with Wayne = 982 + 48*x = 982 + 48x after x months, money with Oscar = 3517 + 9*x = 3517 + 9x since both amounts will be same after x months, 982 + 48x = 3517 + 9x ==> 48x - 9x = 3517 - 982 ==> 39x = 2535 ==> x = 2535/39 = 65 so both Wayne and Oscar will have same amount of money after 65 months i.e 5 years 5 months

OpenStudy (anonymous):

w=982$ and he saves 48 for some months O=3517$ and he saves 9$ for some months let months =x 48x+982=y 3517+9x=y equating both 3517+9x=48x+982 x=65 months

OpenStudy (anonymous):

cool for the guy above me....:) right answer keep it up

OpenStudy (anonymous):

thanks hashsam!!! even u got the number of months right......

OpenStudy (anonymous):

hahaha thanks guys! :D u really helped me a lot

OpenStudy (anonymous):

then give medals to both of us :D

OpenStudy (anonymous):

haha lol right...

OpenStudy (anonymous):

yes, and become our fan also... o( ;})

OpenStudy (anonymous):

lol . what are medals for anyway?

OpenStudy (anonymous):

to show yr appreciation for good help and to mark the question as solved and closed so that other people do not waste time on this question......

OpenStudy (anonymous):

sure :) can u help me with another question later too?

OpenStudy (anonymous):

sure send them across.....